OpenStudy (anonymous):

Which of the following is a solution for the system y = -3x + 5 and y = 2x + 5? (1, 7) (5, 0) (2, -1) (0, 5) I know its not the first 2 but I always make a mistake on these no matter how simple it is, sorry please help

OpenStudy (souvik):

y=-3x+5 and y=2x+5 hence -3x+5=2x+5

OpenStudy (austinl):

Get all the x terms on one side of the equation, and all the numerical terms on the other.

OpenStudy (anonymous):

you lost me?

OpenStudy (souvik):

-3x-2x=5-5 -5x=0 x=0

OpenStudy (anonymous):

o okay I see how you got it thank you

OpenStudy (souvik):

welcome...